# Non-local Hamilton–Jacobi-Bellman equations for the stochastic optimal control of path-dependent piecewise deterministic processes

### v1

@@ 1. Introduction
Piecewise deterministic Markov processes (PDMPs) have been popularized by Davis (see especially [Davis84] and [DavisBook]) and found many applications. PDMPs are pure jump processes; between consecutive jump times, they are deterministic and solve ordinary differential equations. We study a non-Markovian counterpart of PDMPs. More specifically, our processes, which we shall call path-dependent piecewise deterministic processes (PDPs), differ from (and extend) PDMPs inasmuch as all relevant data are allowed to depend on the (realized) process history as opposed to just the (realized) state as is the case for PDMPs. E.g., between consecutive jump times, path-dependent PDPs solve (delay) functional differential equations. Consequently, our theory broadens the realm of possible applications.111 General, not necessarily Markovian, PDPs are briefly mentioned in [Jacobsen]. We are not aware of any thorough treatment of path-dependent PDPs.
Our work focuses on stochastic optimal control problems that involve path-dependent PDPs and their related Hamilton–Jacobi–Bellman (HJB) equations. These equations are non-local and path-dependent. More precisely, our HJB equations are of the following form:
@@
(1.1) $\begin{split}&\partial_{t}u(t,x)+\inf_{a\in A}\Big{\{}\lambda(t,x,a)\,\int_{{\mathbb{R}}^{d}}[u(t,{\mathbf{1}}_{[0,t)}\,x+{\mathbf{1}}_{[t,\infty)}e)-u(t,x)]\,Q(t,x,a,{\mathrm{d}}e)\\ &\qquad+\ell(t,x,a)+(f(t,x,a),\partial_{x}u(t,x))\Big{\}}=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}).\end{split}$
Here, $D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$ denotes the Skorokhod space. Also note that all terms in (1.1) are to be understood as non-anticipating, e.g., $u(t,x)=u(t,x_{\cdot\wedge t})$. Counterparts of (1.1) in the non-path-dependent case would be of the form
@@
(1.2) $\begin{split}&\partial_{t}\tilde{u}(t,\tilde{x})+\inf_{a\in A}\{\tilde{\lambda}(t,\tilde{x},a)\,\int_{{\mathbb{R}}^{d}}[\tilde{u}(t,e)-\tilde{u}(t,\tilde{x})]\,\tilde{Q}(t,\tilde{x},a,{\mathrm{d}}e)\\ &\qquad+\tilde{\ell}(t,\tilde{x},a)+(\tilde{f}(t,\tilde{x},a),\partial_{\tilde{x}}\tilde{u}(t,\tilde{x}))\}=0,\quad(t,\tilde{x})\in[0,T)\times{\mathbb{R}}^{d}.\end{split}$
As a rule, classical solutions to such HJB equations rarely exist and theories of non-smooth solutions such as viscosity or minimax solutions are needed. In the path-dependent case, the only relevant work for non-local equations is [Keller16SPA], where viscosity solutions for semilinear non-local path-dependent partial differential equations are studied. But there have been no works on nonsmooth solutions for fully nonlinear non-local path-dependent partial differential equations such as (1.1). This is in stark contrast to the Markovian case: Partial differential equations associated to PDMPs, i.e., variations of (1.2), were studied in [Vermes85Stoch], [Soner86SICON-II], [Ye94] [DempsterYe95AAP], [DavisFarid], [ForwickEtAl04], [Goreac12], and [BandiniThieullen20AMO] to mention but a few works.
@@ 1.1. A simplified presentation of the control problem
over an appropriate class of controls $\alpha$. The controlled stochastic process $X^{t,x,\alpha}$ is a path-dependent piecewise deterministic process with jump times $T_{n}^{t,x,\alpha}$, $n\in{\mathbb{N}}$, and post-jump locations $E_{n}^{t,x,\alpha}$, $n\in{\mathbb{N}}$, such that the the following holds:
@@
$T_{0}^{t,x,\alpha}=t,\quad\text{$T_{1}^{t,x,\alpha}$ is the first jump time of $X^{t,x,\alpha}$, etc.}$.
@@
${{\mathbb{P}}}(T^{t,x,\alpha}_{1}>s)$ $=F^{t,x,\alpha}(s)\text{ for every $s>T$ and}$
@@
where $\phi^{t,x,\alpha}$ is [^2] Subject to our data satisfy appropriate conditions. (!) the unique solution of $\frac{d}{ds}\phi^{t,x,\alpha}(s)=f(s,\phi_{\cdot\wedge s}^{t,x,\alpha},\alpha(s))$ on $(t,\infty)$ with initial condition $\phi^{t,x,\alpha}(s)=x(s)$ on $[0,t]$, i.e., $\phi^{t,x,\alpha}$ and $X^{t,x,\alpha}$ coincide before the first jump time.
The class of controls: Differences between the Markovian and the path-dependent cases. In both cases, controls are open-loop between jump times. In the Markovian case, admissible controls are of the form [^3] For readability’s sake, we omit the superscript $\phantom{}{}^{t,x,\alpha}$ or part of it.
$\tilde{\alpha}(t)=\tilde{\alpha}_{n}(t|T_{n},E_{n}),\quad T_{n}\leq t<T_{n+1}.$
This structure leads naturally to an equivalent (discrete-time) Markov decision model (see, e.g., [DavisBook]). [^4] Such an equivalency is of major importance in the theory of PDMPs and suitable counterparts are crucial in this work as well. For details, see section 4.5 and appendix A. In the path-dependent case, controls of the form
@@
where $X^{\alpha}$ is our controlled piecewise deterministic process, ought to be at least a subset of all admissible controls. To establish a correspondence to an equivalent discrete-time model, we would need to express the history $\{X^{\alpha}(s)\}_{0\leq s\leq t}$ as an appropriate function of finitely many relevant random variables such as the past jump times $T_{1}$, $\ldots$, $T_{n}$, and past jump locations $E_{1}$, $\ldots$, $E_{n}$. However, to determine $X^{\alpha}$ on $[0,t]$, we additionally [^5] Unlike in the Markovian case, where only $X^{\alpha}(T_{n})=E_{n}$ is needed. need to know the control history, i.e., all (realized) open-loop controls $\alpha_{0}$, $\ldots$, $\alpha_{n-1}$. [^6] This will lead to larger class of admissible controls compared to the ones described in (1.3). What makes matters slightly complicated is that each such (realized) open-loop control $\alpha_{k}$ also depends on $T_{1}$, $\ldots$, $T_{k}$, $E_{1}$, $\ldots$, $E_{k}$, and $\alpha_{0}$, $\ldots$, $\alpha_{k-1}$. To make the situation more tractable, we will use randomized policies (see Definition 4.7). Finally note that such a seemingly complicated setting for our controls is needed in the proof of the dynamic programming principle (Theorem 4.10), which is crucial in establishing that solutions of the HJB equation (1.1) coincide with the value function of our stochastic optimal control problem.
@@
1.2. Our approach for the non-local path-dependent HJB equation (1.1)
@@
We follow the methodology by Davis and Farid in [DavisFarid], where existence and uniqueness of viscosity solutions for HJB equations related to PDMPs are established via a fixed-point argument. However, our situation in the path-dependent case leads to several additional obstacles. We continue by describing some of those difficulties and how they will be circumvented. First note that main ingredients of the Davis–Farid methodology are existence and uniqueness results of certain local HJB equations, in our case, these HJB equations are path-dependent and of the form
(1.4) $\begin{split}\partial_{t}u(t,x)+F_{\psi}(t,x,u(t,x),\partial_{x}u(t,x))=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}),\end{split}$
where $\psi:[0,T]\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\to{\mathbb{R}}$ and $F_{\psi}=F_{\psi}(t,x,y,z)$ is of the form
@@
(1.5) $\begin{split}&F_{\psi}(t,x,y,z)=\inf_{a\in A}\{\lambda(t,x,a)\int_{{\mathbb{R}}^{d}}[\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)-y]\,Q(t,x,a,{\mathrm{d}}e)\\ &\quad+\ell(t,x,a)+(f(t,x,a),z)\},\quad(t,x,y,z)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\times{\mathbb{R}}\times{\mathbb{R}}^{d}.\end{split}$
Roughly speaking, a function $u:[0,T]\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\to{\mathbb{R}}$ is said to be a solution of (1.1) in some sense if $u$ is a solution of (1.4) with $\psi\equiv u$ in the same sense.
@@
One of our obstacles in the path-dependent case is a possible lack of regularity of the term $\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)$ in (1.5). Even if $\psi$ is continuous, the map $t\mapsto\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)$ might be not continuous (see appendix C) and therefore we can expect the Hamiltonian $F_{\psi}$ to be at most measurable with respect to $t$. In our companion paper [BK1], we develop a minimax solution theory for path-dependent Hamilton–Jacobi equations with time-measurable coefficients and also with “$u$-dependence” of the Hamiltonian. There seems to have been no treatment of any of those two aspects in the literature (see [GLP21AMO] for an overview) and thus the results in [BK1] ought to be of independent interest besides the current matter.
@@
A second obstacle is the following. A standard proof of a comparison principle between appropriate sub- and supersolutions of (1.4) (as done in [BK1]) requires at least some regularity of $F_{\psi}$ with respect to its $x$-component. Typically, some type of uniform continuity is required. For simplicity (even though it might be more difficult to prove), let us require Lipschitz continuity, i.e., we would need [^7] We show only the term that causes most trouble.
@@
which, in the Markovian or state-dependent case, is trivially satisfied, as
@@
for some function $\tilde{\psi}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ with $\tilde{\psi}(t,x(t))=\psi(t,x)$. The slightly involved proof of Lemma 5.7 deals with this issue by establishing Lipschitz continuity for our value function $V$, which yields the desired regularity at least for $F_{\psi}$ with $\psi=V$.

### v2

@@ Abstract.
The authors would like to thank the anonymous referees for many valuable suggestions and comments.
@@ 1. Introduction
Since B. Dupire’s introduction of a functional Itô calculus in [dupirefunctional], a significant literature on second order path-dependent partial differential equations (PPDEs) of the form
(1.1) $\partial_{t}u(t,x)+H(t,x,u(t,x),\partial_{x}u(t,x),\partial_{xx}u(t,x))=0,\quad(t,x)\in C({\mathbb{R}}_{+},{\mathbb{R}}^{d}),$
has been developed (see section 1.1 below). Note that the derivatives in (1.1) are to be understood in the sense of [dupirefunctional]. PPDEs are relevant for non-Markovian stochastic optimal control problems, non-Markovian stochastic differential games, and for the pricing of path-dependent options in financial mathematics. Most works deal with problems related to diffusion processes. There are no corresponding theories for fully nonlinear non-local PPDEs, which are relevant for non-Markovian problems related to jump processes. This work aims to fill this gap.
Without our research, already relatively simple classes of problems involving jumps and path-dependence (see Remark 1.2 below) are out of reach. A possible specific application that can be addressed by our theory is presented in section 1.4.
@@
To be more precise, we investigate non-local path-dependent Hamilton–Jacobi–Bellman (HJB) equations that are related to non-Markovian counterparts of piecewise deterministic Markov processes (PDMPs). PDMPs have been popularized by M.H.A. Davis (see especially [Davis84] and [DavisBook]) and found many applications. PDMPs are jump processes with a drift component; between consecutive jump times, they are deterministic and solve ordinary differential equations (ODEs). We call our non-Markovian counterparts path-dependent piecewise deterministic processes (PDPs). They differ from and extend PDMPs inasmuch as all relevant data are allowed to depend on the realized process history as opposed to just the realized state as is the case for PDMPs. E.g., between consecutive jump times, path-dependent PDPs solve delay functional differential equations. Consequently, our theory broadens the realm of possible applications. [^1] General, not necessarily Markovian, PDPs are briefly mentioned on pp. 25–26 in [Jacobsen]. We are not aware of any thorough treatment of path-dependent PDPs.
Our main results are existence and uniqueness for terminal value problems involving non-local path-dependent HJB equations of the form
@@
(1.2) $\begin{split}&\partial_{t}u(t,x)+\inf_{a\in A}\Big\{\lambda(t,x,a)\,\int_{{\mathbb{R}}^{d}}[u(t,{\mathbf{1}}_{[0,t)}\,x+{\mathbf{1}}_{[t,\infty)}e)-u(t,x)]\,Q(t,x,a,{\mathrm{d}}e)\\ &\qquad+\ell(t,x,a)+(f(t,x,a),\partial_{x}u(t,x))\Big\}=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}),\end{split}$
where $D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$ denotes the Skorokhod space and all terms are to be understood as non-anticipating, e.g., $u(t,x)=u(t,x_{\cdot\wedge t})$. Moreover, we show that their solutions coincide with value functions of appropriate stochastic optimal control problems related to path-dependent PDPs. To this end, we establish a suitable dynamic programming principle.
Remark 1.1.
A precise definition of the derivatives in (1.2) is given in subsection 5.1 (note that they are not Fréchet derivatives). Just as in the case of the (local) PPDE (1.1), these derivatives can be understood in the sense of B. Dupire’s so-called horizontal and vertical derivatives in [dupirefunctional]. However, as it is more common in the literature on first order PPDEs, we prefer to use the earlier notion of the so-called coinvariant derivatives on path spaces, which are due to A.V. Kim (see, e.g., [Kim85] and [KimBook]). Both notions of derivatives are essentially equivalent (see, e.g., section 3 in [BK1] for more details).
Counterparts of (1.2) in the non-path-dependent case would be of the form
@@
(1.3) $\begin{split}&\partial_{t}\tilde{u}(t,\tilde{x})+\inf_{a\in A}\Big\{\tilde{\lambda}(t,\tilde{x},a)\,\int_{{\mathbb{R}}^{d}}[\tilde{u}(t,e)-\tilde{u}(t,\tilde{x})]\,\tilde{Q}(t,\tilde{x},a,{\mathrm{d}}e)\\ &\qquad+\tilde{\ell}(t,\tilde{x},a)+(\tilde{f}(t,\tilde{x},a),\partial_{\tilde{x}}\tilde{u}(t,\tilde{x}))\Big\}=0,\quad(t,\tilde{x})\in[0,T)\times{\mathbb{R}}^{d}.\end{split}$
As a rule, classical solutions to such HJB equations rarely exist and theories of non-smooth solutions such as viscosity solutions (see, e.g., [CrandallIshiiLions], [BardiCapuzzoDolcetta]) or minimax solutions (see, e.g., [Subbotin]) are needed. In this work, we use minimax solutions, mainly due to a technical obstacle, which is explained in section 1.3.
Remark 1.2.
Note that our new theory is needed for relatively basic cases.
(i) Optimal control of or option pricing related to standard PDMPs (see, e.g., section 4 in [BR10PDMP] and section 10 in [Jacobsen]) with the same data as the HJB equation (1.3) but with path-dependent terminal cost or path-dependent payoff, e.g., $h(x)=\sup_{0\leq t\leq T}\left|{x(t)}\right|$. Then the corresponding HJB equation is a PPDE of the form
$\partial_{t}u(t,x)+\inf_{a\in A}\{\tilde{\lambda}(t,x(t),a)\,\int_{{\mathbb{R}}^{d}}[u(t,{\mathbf{1}}_{[0,t)}\,x+{\mathbf{1}}_{[t,\infty)}e)-u(t,x)]\,\tilde{Q}(t,x(t),a,{\mathrm{d}}e)$
$\qquad+\tilde{\ell}(t,x(t),a)+(\tilde{f}(t,x(t),a),\partial_{x}\tilde{u}(t,x))\}=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}),$
with terminal condition $u(T,x)=h(x)$, $x\in D({\mathbb{R}}_{+},{\mathbb{R}}^{d})$. Note that $A$ would be a singleton in the case of option pricing.
(ii) Change of the ODE in a PDMP to a delay differential equation. Even results on well-posedness for HJB equations associated to the optimal control of delay differential equations with jumps that follow a Poisson process are new. If $d=1$, then the corresponding HJB equation would be of the form
$\partial_{t}u(t,x)+\lambda[u(t,{\mathbf{1}}_{[0,t)}\,x+{\mathbf{1}}_{[t,\infty)}\,(x(t)+1))-u(t,x)]$
$\qquad+\inf_{a\in A}(\tilde{f}(t,x((t-\tau)\vee 0),a),\partial_{x}u(t,x))=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}),$
as a special case of (1.2). Similar and actually more general models with jumps and delays are used in [AH20] for option pricing.
1.1. Related research
PDEs associated to PDMPs, i.e., variations of (1.3), were studied in [Vermes85Stoch], [Soner86SICON-II], [Ye94] [DempsterYe95AAP], [DavisFarid], [ForwickEtAl04], [Goreac12], [Bandini18COCV], [Bandini19SICON], [ACE19SIFIN], [BandiniThieullen20AMO], and [ColaneriEtAl20SPA], to mention but a few works.
Remark 1.3.
The data in the above literature are either not time-dependent or required to be continuous with respect to time. Our results are stronger in the sense that we require our data to be only measurable with respect to time.
Next, we provide a short overview of the literature on PPDEs.
Existence and uniqueness of minimax solutions for first order path-dependent Hamilton–Jacobi equations was established in [Lukoyanov03] and of viscosity solutions in [Lukoyanov07]. Wellposedness of viscosity solutions for first order path-dependent HJB equations under significantly weaker conditions (only continuity of the data in the path variable $x$ with respect to sup norm is needed instead of requiring continuity with respect to an $L_{p}$ norm) was proved in [zhou2020viscosity1st]. For more details as well as a comprehensive overview for first order PPDEs, we refer to the recent survey paper [GL24survey].
Wellposedness of viscosity solutions for semilinear second order PPDEs was established in [EKTZ11] and, subsequently, for semilinear second order path-dependent integro-differential equations in [Keller16SPA]. Note that [Keller16SPA] is the only relevant work for non-local PPDEs and that there have been no works on nonsmooth solutions for fully nonlinear non-local PPDEs such as (1.2).
Many works in the literature on fully nonlinear second order PPDEs such as [ETZ_I], [ETZ_II], [EkrenZhang16PUQR], [RTZ17], [RR20SIMA], [cosso21v2path], [zhou2020viscosity], and [Keller24] cover HJB equations of the form
(1.4) $\begin{split}&\partial_{t}u(t,x)+\inf_{a\in A}\left\{\frac{1}{2}\sigma(t,x,a)^{2}\partial_{xx}u(t,x)+\ell(t,x,a)+f(t,x,a)\partial_{x}u(t,x)\right\}=0,\\ &\qquad\qquad(t,x)\in[0,T)\times C({\mathbb{R}}_{+},{\mathbb{R}}^{d}).\end{split}$
Regarding existence and uniqueness of viscosity solutions for second order path-dependent HJB equations such as (1.4), [zhou2020viscosity] and [Keller24] allow for the weakest conditions [^2] We refer to subsection 1.3 in [Keller24] for a more detailed discussion. among the just mentioned works such as a fairly general possibly degenerated controlled diffusion term as well as continuity of the data in the sup norm (as opposed to continuity in an $L_{p}$-norm). It should be mentioned that the notion of viscosity solution in [zhou2020viscosity] is a modification of the usual one in the non-path-dependent case (see [CrandallLions83TAMS]) and uses pointwise tangency for the employed test functions whereas the notion in [Keller24] uses tangency in mean (which was introduced in [EKTZ11]).
Remark 1.4.
Instead of the controlled diffusion term $(1/2)\sigma(t,x,a)^{2}\partial_{xx}u(t,x)$ in (1.4), we have the comparable controlled non-local term
$\int_{{\mathbb{R}}^{d}}[u(t,{\mathbf{1}}_{[0,t)}\,x+{\mathbf{1}}_{[t,\infty)}e)-u(t,x)]\lambda(t,x,a)\,Q(t,x,a,{\mathrm{d}}e)$
in (1.2). Our work allows for the data to be continuous in $x$ with respect to the sup norm as well. Furthermore, our data need only be measurable with respect to time whereas in [zhou2020viscosity] continuity in time is required.
Last but not least, we want to note that, besides the PPDE approach (based on path derivatives by B. Dupire [dupirefunctional] or A.V. Kim [KimBook]), it is also very common for the treatment of stochastic optimal control problems with delays to use the Hilbert space $L^{2}(-\tau,0;{\mathbb{R}}^{d})$ or the space $C([-\tau,0],{\mathbb{R}}^{d})$ as state space and consider HJB equations on those spaces with derivatives understood as Fréchet derivatives: For a corresponding overview, we refer the reader to the monograph [FabbriGozziSwiech] (in particular to its section 2.6.8 and the references therein).
@@ 1.2. A simplified presentation of the control problem
over an appropriate class of controls $\alpha$, which will be discussed shortly. The controlled stochastic process $X^{t,x,\alpha}$ is a path-dependent PDP with jump times $T_{n}^{t,x,\alpha}$, $n\in{\mathbb{N}}$, and post-jump locations $E_{n}^{t,x,\alpha}$, $n\in{\mathbb{N}}$, such that the the following holds:
@@
[…]
@@
${{\mathbb{P}}}(T^{t,x,\alpha}_{1}>s)$ $=F^{t,x,\alpha}(s)\text{ for every $s>{\color[rgb]{0,0,0}t}$ and}$
@@
where $\phi^{t,x,\alpha}$ is the unique solution of $\frac{d}{ds}\phi^{t,x,\alpha}(s)=f(s,\phi_{\cdot\wedge s}^{t,x,\alpha},\alpha(s))$ on $(t,\infty)$ with initial condition $\phi^{t,x,\alpha}(s)=x(s)$ on $[0,t]$, i.e., $\phi^{t,x,\alpha}$ and $X^{t,x,\alpha}$ coincide before the first jump time.
The class of controls: Differences between the Markovian and the path-dependent cases. In both cases, controls are open-loop between jump times.
In the Markovian case, admissible controls are of the form [^3] For readability’s sake, we omit the superscript $\phantom{}{}^{t,x,\alpha}$ or part of it.
(1.5) $\tilde{\alpha}(t)=\tilde{\alpha}_{n}(t|T_{n},E_{n}),\quad T_{n}\leq t<T_{n+1}.$
This structure leads naturally to an equivalent (discrete-time) Markov decision model (see, e.g., [DavisBook]). Such an equivalency is of major importance in the theory of PDMPs and a suitable counterpart is crucial in this work as well.
In the path-dependent case, our controls are of the form
@@
where $X^{\alpha}$ is our controlled piecewise deterministic process. In the spirit of the Markovian case, we establish a correspondence between our continuous-time control problem and a related discrete-time model (see section 4.4 and especially Theorem 4.15). To this end, we express the history $\{X^{\alpha}(s)\}_{0\leq s\leq t}$ in (1.6) as an appropriate function that depends on the past jump times $T_{1}$, $\ldots$, $T_{n}$, the past jump locations $E_{1}$, $\ldots$, $E_{n}$, and, in contrast to the Markovian case, also on the control history, i.e., on all realized open-loop controls (see Definition 4.9 for the precise definition of the policies for our control problem). Doing so is needed for the proof of the dynamic programming principle (Theorem 4.17), which is crucial in establishing that solutions of the HJB equation (1.2) coincide with the value function of our stochastic optimal control problem.
@@
1.3. Our approach in dealing with specific difficulties concerning the non-local path-dependent HJB equation (1.2)
@@
We follow the methodology by Davis and Farid in [DavisFarid], where existence and uniqueness of viscosity solutions for HJB equations related to PDMPs are established via a fixed-point argument (see the beginning of section 5 for a detailed outline). However, our situation in the path-dependent case leads to several additional obstacles. We continue by describing some of those difficulties and how they will be circumvented. First note that the main ingredients of the Davis–Farid methodology are existence and uniqueness results of certain local HJB equations, in our case, these HJB equations are path-dependent and of the form
(1.7) $\begin{split}\partial_{t}u(t,x)+{\color[rgb]{0,0,0}H^{\psi}(t,x,u(t,x),\partial_{x}u(t,x))}=0,\quad(t,x)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d}),\end{split}$
where $\psi:[0,T]\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\to{\mathbb{R}}$ and $\color[rgb]{0,0,0}H^{\psi}(t,x,y,z)$} is of the form
@@
(1.8) $\begin{split}&{\color[rgb]{0,0,0}H^{\psi}(t,x,y,z)}=\inf_{a\in A}\{\lambda(t,x,a)\int_{{\mathbb{R}}^{d}}[\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)-y]\,Q(t,x,a,{\mathrm{d}}e)\\ &\quad+\ell(t,x,a)+(f(t,x,a),z)\},\quad(t,x,y,z)\in[0,T)\times D({\mathbb{R}}_{+},{\mathbb{R}}^{d})\times{\mathbb{R}}\times{\mathbb{R}}^{d}.\end{split}$
Note that (1.2) is identical to (1.7) in case $\psi\equiv u$.
@@
One of our obstacles in the path-dependent case is a possible lack of regularity of the term $\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)$ in (1.8). Even if $\psi$ is continuous, the map $t\mapsto\psi(t,{\mathbf{1}}_{[0,t)}x+{\mathbf{1}}_{[t,\infty)}e)$ might not be continuous (see appendix C) and therefore we can expect the Hamiltonian $F_{\psi}$ to be at most measurable with respect to $t$. To deal with this issue, we develop in our companion paper [BK1] a minimax solution theory for path-dependent Hamilton–Jacobi equations with time-measurable coefficients and also with “$u$-dependence” of the Hamiltonian, i.e., we cover equations of the form $\partial_{t}u+H(t,x,u,\partial_{x}u)=0$, whereas the previous minimax solutions literature on PPDEs covers equations of the form $\partial_{t}u+H(t,x,\partial_{x}u)=0$. There seems to have been no treatment of any of those two aspects in the literature (see [GLP21AMO] for an overview) and thus the results in [BK1] ought to be of independent interest besides the current matter. Also note that, while we use in [BK1] a combination of minimax and viscosity solution techniques to establish uniqueness for minimax solutions and to verify that value functions of relevant optimal control problems are minimax solutions of the corresponding HJB equations, a complete viscosity solution theory for PPDEs with time-measurable coefficients does not exist to the best of our knowledge. For further discussion on relevant background and difficulties, we refer to section 1 in [BK1]. Furthermore, note that there has been no minimax solution theory related to optimal control of PDMPs.
@@
A second obstacle is the following. A standard proof of a comparison principle between appropriate sub- and supersolutions of (1.7) (as done in [BK1]) requires at least some regularity of $\color[rgb]{0,0,0}H^{\psi}$} with respect to its $x$-component. Typically, some type of uniform continuity is required. For simplicity (even though it might be more difficult to prove), let us require Lipschitz continuity, i.e., we would need
@@
which, in the Markovian or state-dependent case, is trivially satisfied because
@@
for some function $\tilde{\psi}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ with $\tilde{\psi}(t,x(t))=\psi(t,x)$. The slightly involved proof of Theorem 5.12 deals with this issue by establishing Lipschitz continuity for our value function $V$, which yields the desired regularity at least for $\color[rgb]{0,0,0}H^{\psi}$} with $\psi=V$.
1.4. The delayed Hodgkin–Huxley model as possible application
The Hodgkin–Huxley (HH) model, developed in 1952 by A. Hodgkin and A. Huxley in [HH52], describes how action potentials (nerve impulses) are initiated and propagated along axons. It is based on the concept of voltage-gated ion channels that allow ions to flow through the membrane of the neuron, generating electrical signals in response to changes in the membrane potential. In the original HH model, the transitions between different states of ion channels (e.g., from closed to open or inactivated states) are deterministic and described by ODEs with specific kinetic parameters. In reality, the behaviour of the ion channels is subject to random fluctuations, and it is natural to model the state transitions through a continuous-time Markov chain $d_{t}$, see e.g., [Kampen].
For such a model one can consider optogenetics control problems. Optogenetics allows precise control of cellular activity by using light to manipulate genetically engineered ion channels that are sensitive to light. In the control problem new (rhodopsin) channels that are sensitive to light are inserted in the neuron. Experimentally, the channel is illuminated and the effect of the illumination is to put the channel in one of its conductive states. From a mathematical point of view, this type of problems has been successfully studied (in the case of spatio-temporal models) by dealing with the theory of controlled infinite dimensional PDMPs, see e.g., [BandiniThieullen20AMO], [RTT17PDMP].
On the other hand, recent studies revealed that significant delays could take place between voltage changes and ion channel activation/inactivation dynamics. As a matter of fact, the response of the gating variables to changes in the membrane potential is not instantaneous, but rather involves some time lag: ion channels may have intrinsic delays in their opening and closing, as well as delays in the response to changes in membrane voltage. The need to capture more realistic dynamics of neurons have lead to the development of the so-called delayed HH models, see e.g. [CessacVolkov]. In this extended model, the membrane potential $v$ would follow a delayed ODE of the type
$C_{m}v^{\prime}(t)=-I_{ion}(v(t),m(t-\tau))+I_{ext}(t),\quad v(0)=v\in[0,V_{max}].$
Here $I_{ion}(v(t),d(t-\tau))$, that represents the ionic current, would depend on the delayed state of the gating variable $d(t-\tau)$, and the stochastic gating state $d$ would be a pure jump process evolving according to a jump measure with compensator $\lambda(v(t-\tau),d(t-\tau))Q(v(t-\tau),d(t-\tau),\rm dy)\,{\color[rgb]{0,0,0}{\mathrm{d}}t}$.
Our theory is well suited to manage such a model. Therefore, as a future application, corresponding optogenetic control problems could be studied by suitably adapting our results to the bounded domain framework.